The following conversion from y-intercept to standard form is incorrect. Explain where the error(s) occurs and how to correct the mistake(s)
Y=1/4x+5
4(y)=4(1/4)x+5
4y=1x+5
4y-x=x-x+5
4y-x=5
4y-x-5=5-5
-x+4y-5=0

Found 2 solutions by jim_thompson5910, josgarithmetic:
Answer by jim_thompson5910(35256)

The mistake happens in line 2. The person doesn't multiply both sides by 4 properly. Note how the 4 on the right side multiplies with only the 1/4x term and not with the 5 also.

Instead, you should distribute the outer 4 to each term inside the parenthesis like shown here
4(1/4x + 5) = 4(1/4x) + 4(5) = x + 20

Therefore,
y = (1/4)x + 5
becomes
4y = x + 20
after you multiply both sides by 4 and distribute

From here, we follow similar steps as shown in the given problem to get the equation into Ax+By+C = 0 form

4y = x + 20
4y - x = x + 20 - x ... subtract x from both sides
-x + 4y = 20
-x + 4y - 20 = 20-20 .... subtract 20 from both sides
-x + 4y - 20 = 0

Answer by josgarithmetic(39617)

Already been answered here:
https://www.algebra.com/algebra/homework/Linear-equations/Linear-equations.faq.question.1133297.html
https://www.algebra.com/algebra/homework/Linear-equations/Linear-equations.faq.question.1133297.html